# Graph Neural Networks in EEG-based Emotion Recognition: A Survey

### v5

@@ 5 Graph-level Stage
There are four graph structures included in the Graph-level Stage: Multi-graph, Hierarchical graph, Time series graph, and Sparse graph. According to the dependencies that graphs indicate, Multi-graph is divided into Horizontal&Vertical graph, Temporal&Frequency graph, and Local&Globle graph. Hierarchical graph is divided into Dynamic hierarchy graph and Predetermined hierarchy graph. The realisation of Time series graph is divided into Temporal graph and Temporal encoder. The sparsity implementation of Sparse graph consists of Threshold and Sparse weight.
@@ 5.1 Multi-graph
Multi-graph allows the model to capture different emotional EEG dependencies simultaneously by concatenating multiple types of graph embedding. Various relational dependencies exist in emotional EEG, such as temporal dependency, frequency dependency, local brain region dependency, etc. The multi-graph structure can learn and consider different dependencies simultaneously, which helps to build relationships in emotional EEG more comprehensively. However, the multi-stream graph modeling process increases the computational burden. The existing Multi-graph can be subdivided into three categories, as shown in Table. 4.
Horizontal&Vertical graph captures both vertical and horizontal spatial dependencies of brain regions. There are differences in the horizontal and vertical connectivity patterns of brain regions so that the Horizon&Vertical structure can model more physiologically plausible spatial dependency. ERHGCN Zheng et al. (2021) and BiHDM Li et al. (2020) both utilize this graph. BiHDM further constructed different horizontal and vertical graphs in the left and right hemispheres based on the existence of imbalance between hemispheres.
Temporal&Frequency graph combines the temporal dependency and frequency dependency of the emotional EEG signal. Temporal dependence allows the model to notice amplitude features that are highly correlated with emotion. Frequency dependence encompasses activating the corresponding frequency band in a particular emotion, such as the increase in the $\alpha$ band during the positive emotion. Existing methods MD-AGCN Li et al. (2021b), HetEmotionNet Jia et al. (2021), and DBGC-AFFNet-AFTL Sun et al. (2022) all adopt a similar two-stream structure, where each stream corresponds to the temporal-spatial and frequency-spatial space, and finally graph embedding are concatenated and fed into the classifier.
Local&Globle graph combines global dependency between brain regions and local dependency within brain regions. In emotional states, global dependence responds to interactions between brain regions, such as activity in surrounding brain regions caused by activity in specific brain regions. Local dependence focuses on relationships within specific brain regions, e.g., activity within the frontal lobe and cortex is associated with positive emotions. GECNN Song et al. (2021b), HD-GCNYe et al. (2022), AHGCNXue et al. (2022), and LGGNetDing et al. (2023) all infer the spatial dependencies of all channels and channels within specific brain regions. Similarly, MRGCN Qiu et al. (2023) introduces short-range and long-range spatial dependencies corresponding to localized intra-region correlations and inter-region correlations.
@@
Graph
Baseline
@@
BiHDM Li et al. (2020)
ERHGCN Zheng et al. (2021)
@@
MD-AGCN Li et al. (2021b)
HetEmotionNet Jia et al. (2021)
DBGC-ATFFNet-AFTL Sun et al. (2022)
@@
GECNN Song et al. (2021b)
HD-GCN Ye et al. (2022)
AHGCN Xue et al. (2022)
MRGCN Qiu et al. (2023)
LGGNet Ding et al. (2023)
5.2 Hierachical graph
@@
Hierarchical graph groups channels to allow the model to capture local spatial dependency within specific brain regions. The spatial dependency of emotional EEG consists of inter-region and intra-region dependencies. The former corresponds to the region-level synergy between different brain regions under emotion, and the latter corresponds to the neural-level activity within the brain region associated with a specific emotion, e.g., the frontal cortex and lobe are related to positive emotions. Therefore, the hierarchical graph structure allows the model to effectively utilize the signal features contained in these local brain regions that are directly related to emotions. The difference with the Local&Globle multi-graph structure is that the hierarchical graph structure does not necessarily employ a multi-stream process, thus reducing the computational burden. The existing Hierarchical graph contains two categories, as shown in Table. 5.
@@
Dynamic hierarchy graph indicates that the model adaptively groups channels and then infers spatial dependency between groups. There are significant individual differences in emotional EEG between subjects, which implies that different subjects will have subtle differences in brain region correlations during the same emotion. Therefore, dynamic hierarchy can reduce the effect of individual differences. AHGCN Xue et al. (2022), SCC-MPGCN Zhao et al. (2022), and MDGCN-SRCNN Bao et al. (2022) firstly infer channel-level spatial dependency, then implement a trainable weight assignment matrix to assign different electrodes to different brain regions, and finally build region-level spatial graph. MST-GNN Liu et al. (2022a) also implements channel-level dependency first, and then, based on the minimum spanning tree, it starts to search for children nodes with the strongest correlation so that each branch can be regarded as a group.
@@
Predetermined hierarchy graph denotes the setting of the brain regions to which the electrodes belong based on existing prior knowledge. Prior knowledge refers to the cytoarchitecture of the brain, i.e., the division of brain regions according to the function. Predetermined hierarchy graph is physiologically reasonable and does not introduce additional computational burden. GMSS Li et al. (2022b), R2G-STNN Li et al. (2019), and VPR Zhang et al. (2020) apply a local-to-global process. They divide electrodes into multiple regions in advance based on prior knowledge and first infer spatial dependency within regions and then between regions. In contrast, V-IAG Song et al. (2021a) utilizes a global-to-local process. It first infers fully connected spatial dependency and then divides electrode regions and aggregates region-level nodes to construct relationships between regions. HD-GCN Ye et al. (2022), and LGGNet Ding et al. (2023) adopt the multi-graph structure in Section. 5.1, where local and global dependencies are reasoned in parallel and then concatenated together.
@@
Graph
Baseline
@@
AHGCN Xue et al. (2022)
SCC-MPGCN Zhao et al. (2022)
MST-GNN Liu et al. (2022a)
MDGCN-SRCNN Bao et al. (2022)
@@
R2G-STNNLi et al. (2019)
VPR Zhang et al. (2020)
V-IAG Song et al. (2021a)
HD-GCN Ye et al. (2022)
GMSS Li et al. (2022b)
LGGNet Ding et al. (2023)
@@ 5.3 Time series graph
Time series graph indicates that the model decomposes the signal into multiple time slices and constructs temporal dependency between slices. The temporal dependency of emotional EEG directly reflects the relationship between signal amplitude change and emotion. A high degree of activation within 300ms in the frontal cortex channels is usually associated with positive emotions, while a low degree of activation is usually associated with negative emotions. Therefore, Time series graph enables the model to construct and exploit the spatial and temporal dependencies of emotional EEG. Typically, it is implemented by splitting the input signal into multiple time slices to model spatial graphs within slices and then deriving the temporal dependency between slices by temporal graph or temporal encoder, as shown in Table. 6.
Temporal graph represents the model regards spatial graph embedding of time slices as nodes to construct a temporal graph. CR-GAT Liu et al. (2022b) and GIGN Ding and Guan (2023) employ a GNN in the temporal dimension to construct a temporal graph by regarding time slices as nodes. MD-AGCN Li et al. (2021b) couples the spatial graphs of all time slices and then takes the average as a temporal graph.
Temporal encoder represents that the model combines GNN and temporal encoder to derive the spatial and temporal dependencies of the emotional EEG. This derivation process is usually divided into two steps; firstly, the GNN is used to infer the spatial dependency within the time slice and update the spatial graph embedding; then, the graph embedding of all the time slices is concatenated together and fed into a temporal encoder. ST-GCLSTM Feng et al. (2022), ASTG-LSTM Li et al. (2021c) and R2G-STNN Li et al. (2019) use LSTM as the temporal encoder. Similarly, HetEmotionNet Jia et al. (2021) uses GRUs, and EmoGT Jiang et al. (2023) utilizes a transformer encoder.
@@
Methods
Baseline
@@ Methods
MD-AGCN Li et al. (2021b)
CR-GAT Liu et al. (2022b)
GIGN Ding and Guan (2023)
@@
R2G-STNN Li et al. (2019)
ASTG-LSTM Li et al. (2021c)
HetEmotionNet Jia et al. (2021)
ST-GCLSTM Feng et al. (2022)
EmoGTJiang et al. (2023)
@@ 5.4 Sparse graph
Sparse graph implies that the graph connectivity relations are sparse. There is a concentration of brain activity in areas directly related to emotion, which means that most connections are weakly associated with emotion. Therefore, specific emotions can be identified by the electrical signals that are active around specific brain regions, such as the frontal cortex and lobe. The sparse graph structure reduces the influence of irrelevant information and retains only the essential connections related to a specific emotion. It differs from the hierarchical structure in that it couples specific groups of channels to achieve fewer connections, whereas the sparse graph structure retains fewer edges by filtering directly on top of the fully connected graph. Existing methods set Threshold or add Sparse weights to the loss function to retain a specific number of edges to form a sparse graph. DAGAM Xu et al. (2023) and SOGNN Li et al. (2021a) set Threshold $k$ to achieve sparsity. DAGAM proposes self-attention graph pooling that regards the Score matrix as the sparse weight matrix and retrains top-$k$ elements to retrain top score edges. SOGNN utilizes a $1\times 2$ max pooling layer to retrain top-$k$ edges. SGA-LSTM Liu et al. (2019), and SparseDGCNN Zhang et al. (2021) set Sparse weight to measure the contributions of different channels. The sparse weight is added to the loss to update the edge matrix. The sparsity is achieved when some elements of the edge matrix are close to 0.
@@ 6 Future Directions
In this paper, we categorize existing methods from the perspective of three stages of constructing GNNs in EEG-based emotion recognition. Despite the progress made by existing methods, there are still some challenges and possible future directions. We analyze several directions that are significant for future EEG-based emotion recognition.
Temporal full-connected graph represents a fully-connected temporal graph across time slices. It aims to solve the problem of incomplete temporal dependency in current EEG-based emotion recognition. The temporal dependency of emotional EEG inferred by existing GNNs lacks the correlation between different channels across time slices. For Time series graph mentioned in the Section. 5.3, although the relationship between time slices can be established by the Temporal graph or encoder, this relationship is limited between the same electrodes across time slices, such as $v_{i}^{t-1}$ and $v_{i}^{t}$, while relations between different electrodes across time slice are ignored, such as $v_{i}^{t-1}$ and $v_{j}^{t}$. This hetero-electrode relationship across time slice corresponds to a delayed response in the emotional state of the brain, i.e., there is asynchrony in the interaction between brain regions. The existence of delays between specific brain regions and surrounding brain regions for brain region activity elicited by emotional excitation implies that this asynchronous relationship contains information directly related to emotion. Therefore, a more complete temporal dependency constructed by the Temporal fully-connected graph is expected to improve the accuracy of emotion recognition.
Graph condensation aims to compress a large and complex graph structure into a simple and compact representation and retain essential emotion-related information. Both coarsening and sparsifying are promising techniques for future EEG-based emotion recognition. Emotional EEG tends to move towards high resolution. Current GNNs in EEG-based emotion recognition utilize fully connected graphs, where nodes usually have an excessive number of neighbors, which provides redundant information and even noise. As the number of electrodes increases, this defect becomes more significant. Therefore, graph condensation techniques can not only improve training efficiency but also weaken the influence of emotionally irrelevant information. The activity of brain regions in emotional states is focused, which implies that the information of specific brain regions is sufficient to identify specific emotions, while most of the connections are emotionally irrelevant. In practice, GNNs in EEG-based emotion recognition are trained to preserve localization. Existing GNNs in this field generally need to run the training process at least once to find the optimal graph condensation, so if the optimal graph is identified earlier, e.g., by training for only a few epochs, then this will significantly increase the running process of GNNs to build emotional representation.
@@
[…]

### v6

@@ 5 Graph-level Stage
There are four graph structures included in the Graph-level Stage: Multi-graph, Hierarchical graph, Time series graph, and Sparse graph. According to the dependencies that graphs indicate, Multi-graph is divided into Horizontal&Vertical graph, Temporal&Frequency graph, and Local&Global graph. Hierarchical graph is divided into Dynamic hierarchy graph and Predetermined hierarchy graph. The realisation of Time series graph is divided into Temporal graph and Temporal encoder. The sparsity implementation of Sparse graph consists of Threshold and Sparse weight.
@@ 5.1 Multi-graph
Multi-graph allows the model to capture different emotional EEG dependencies simultaneously by concatenating multiple types of graph embeddings. Various relational dependencies exist in emotional EEG, such as temporal, frequency, and local brain region dependency, etc. The multi-graph structure can learn and consider different dependencies simultaneously, which helps to build relationships in emotional EEG more comprehensively. However, the multi-stream graph modeling process increases the computational burden. The existing Multi-graph can be subdivided into three categories, as shown in Table 4.
Horizontal&Vertical graph captures both vertical and horizontal spatial dependencies of brain regions. There are differences in the horizontal and vertical connectivity patterns of brain regions so that the Horizontal&Vertical structure can model more physiologically plausible spatial dependency. ERHGCN [45] and BiHDM [17] both utilize this graph. BiHDM further constructed different horizontal and vertical graphs in the left and right hemispheres based on the existence of imbalance between hemispheres.
Temporal&Frequency graph combines the temporal dependency and frequency dependency of the emotional EEG signal. Temporal dependence allows the model to notice amplitude features that are highly correlated with emotion. Frequency dependence encompasses activating the corresponding frequency band in a particular emotion, such as the increase in the $\alpha$ band during the positive emotion. Existing methods MD-AGCN [13], HetEmotionNet [8], and DBGC-ATFFNet-AFTL [34] all adopt a similar two-stream structure, where each stream corresponds to the temporal-spatial and frequency-spatial space, and finally graph embeddings are concatenated and fed into the classifier.
Local&Global graph combines global dependency between brain regions and local dependency within brain regions. In emotional states, global dependence responds to interactions between brain regions, such as activity in surrounding brain regions caused by activity in specific brain regions. Local dependence focuses on relationships within specific brain regions, e.g., activity within the frontal lobe and cortex is associated with positive emotions. GECNN [32], HD-GCN [38], AHGCN [36], and LGGNet [4] all infer the spatial dependencies of all channels and channels within specific brain regions. Similarly, MRGCN [29] introduces short-range and long-range spatial dependencies corresponding to localized intra-region correlations and inter-region correlations.
@@
Graph Baseline Graph Details
@@
BiHDM [17]
ERHGCN [45]
Bi-hemispheric Paired Discrepancy
Brain Regions
@@
MD-AGCN [13]
HetEmotionNet [8]
DBGC-ATFFNet-AFTL [34]
VBH-GNN [21]
Two-stream Structure
Heterogeneous Domain Fusion
Two-stream Structure
Two-stream Structure
@@
GECNN [32]
HD-GCN [38]
AHGCN [36]
MRGCN [29]
LGGNet [4]
Electrodes Groups
Brain Region
Brain Region
Long/Short Range Dependencies
Brain Region
5.2 Hierarchical graph
@@
Hierarchical graph groups channels to allow the model to capture local spatial dependency within specific brain regions. The spatial dependency of emotional EEG consists of inter-region and intra-region dependencies. The former corresponds to the region-level synergy between different brain regions under emotion, and the latter corresponds to the neural-level activity within the brain region associated with a specific emotion, e.g., the frontal cortex and lobe are related to positive emotions. Therefore, the hierarchical graph structure allows the model to effectively utilize the signal features contained in these local brain regions that are directly related to emotions. The difference from the Local&Global multi-graph structure is that the hierarchical graph structure does not necessarily employ a multi-stream process, thus reducing the computational burden. The existing Hierarchical graph contains two categories, as shown in Table 5.
@@
Dynamic hierarchy graph indicates that the model adaptively groups channels and then infers spatial dependency between groups. There are significant individual differences in emotional EEG between subjects, which implies that different subjects will have subtle differences in brain region correlations during the same emotion. Therefore, dynamic hierarchy can reduce the effect of individual differences. AHGCN [36], SCC-MPGCN [43], and MDGCN-SRCNN [1] firstly infer channel-level spatial dependency, then implement a trainable weight assignment matrix to assign different electrodes to different brain regions, and finally build region-level spatial graph. MST-GNN [22] also implements channel-level dependency first, and then, based on the minimum spanning tree, it starts to search for children nodes with the strongest correlation so that each branch can be regarded as a group.
@@
Predetermined hierarchy graph denotes the setting of the brain regions to which the electrodes belong based on existing prior knowledge. Prior knowledge refers to the cytoarchitecture of the brain, i.e., the division of brain regions according to the function. Predetermined hierarchy graph is physiologically reasonable and does not introduce additional computational burden. GMSS [16], R2G-STNN [18], and VPR [41] apply a local-to-global process. They divide electrodes into multiple regions in advance based on prior knowledge and first infer spatial dependency within regions and then between regions. In contrast, V-IAG [31] utilizes a global-to-local process. It first infers fully connected spatial dependency and then divides electrode regions and aggregates region-level nodes to construct relationships between regions. HD-GCN [38], and LGGNet [4] adopt the multi-graph structure in Section 5.1, where local and global dependencies are reasoned in parallel and then concatenated together.
@@
Graph Type Baseline Hierarchy Details
@@
AHGCN [36]
SCC-MPGCN [43]
MST-GNN [22]
MDGCN-SRCNN [1]
Channels/Regions Spatial Dependency
Channels/Regions Spatial Dependency
Channels/Regions Spatial Dependency
Correlation-based Tree Structure
@@
R2G-STNN [18]
VPR [41]
V-IAG [31]
HD-GCN [38]
GMSS [16]
LGGNet [4]
Local to Global
Local to Global
Global to Local
Multi-Graph Concatenation
Local to Global
Multi-Graph Concatenation
@@ 5.3 Time series graph
Time series graph indicates that the model decomposes the signal into multiple time slices and constructs temporal dependency between slices. The temporal dependency of emotional EEG directly reflects the relationship between signal amplitude change and emotion. A high degree of activation within 300ms in the frontal cortex channels is usually associated with positive emotions, while a low degree of activation is usually associated with negative emotions. Therefore, Time series graph enables the model to construct and exploit the spatial and temporal dependencies of emotional EEG. Typically, it is implemented by splitting the input signal into multiple time slices to model spatial graphs within slices and then deriving the temporal dependency between slices by temporal graph or temporal encoder, as shown in Table 6.
Temporal graph represents the model regards spatial graph embeddings of time slices as nodes to construct a temporal graph. CR-GAT [23], VSGT [20] and GIGN [3] employ a GNN in the temporal dimension to construct a temporal graph by regarding time slices as nodes. MD-AGCN [13] couples the spatial graphs of all time slices and then takes the average as a temporal graph.
Temporal encoder represents that the model combines GNN and temporal encoder to derive the spatial and temporal dependencies of the emotional EEG. This derivation process is usually divided into two steps; firstly, the GNN is used to infer the spatial dependency within the time slices and update the spatial graph embeddings; then, the graph embedding of all the time slices is concatenated together and fed into a temporal encoder. ST-GCLSTM [7], ASTG-LSTM [15], R2G-STNN [18] and DTC-GCN [49] use LSTM as the temporal encoder. Similarly, HetEmotionNet [8] uses GRUs, and EmoGT [9] and EmT [5] utilizes a transformer encoder.
@@
Methods Baseline Temporal Details
@@
MD-AGCN [13]
CR-GAT [23]
GIGN [3]
VSGT [20]
Average Spatial Graphs
Time Slices as Nodes
Time Slices as Nodes
Time Slices as Nodes
@@
R2G-STNN [18]
ASTG-LSTM [15]
HetEmotionNet [8]
ST-GCLSTM [7]
EmoGT [9]
EmT [5]
DTC-GCN [49]
LSTM
LSTM
GRUs
LSTM
Transformers
Transformers
LSTM
@@ 5.4 Sparse graph
Sparse graph implies that the graph connectivity relations are sparse. There is a concentration of brain activity in areas directly related to emotion, which means that most connections are weakly associated with emotion. Therefore, specific emotions can be identified by the electrical signals that are active around specific brain regions, such as the frontal cortex and lobe. The sparse graph structure reduces the influence of irrelevant information and retains only the essential connections related to a specific emotion. It differs from the hierarchical structure in that it couples specific groups of channels to achieve fewer connections, whereas the sparse graph structure retains fewer edges by filtering directly on top of the fully connected graph. Existing methods set Threshold or add Sparse weights to the loss function to retain a specific number of edges to form a sparse graph. DAGAM [35] and SOGNN [12] set Threshold $k$ to achieve sparsity. DAGAM proposes self-attention graph pooling that regards the Score matrix as the sparse weight matrix and retains top-$k$ elements to retain top score edges. SOGNN utilizes a $1\times 2$ max pooling layer to retain top-$k$ edges. SGA-LSTM [24], and SparseDGCNN [40] set Sparse weight to measure the contributions of different channels. The sparse weight is added to the loss to update the edge matrix. The sparsity is achieved when some elements of the edge matrix are close to 0.
@@ 6 Future Directions
Despite the progress made by existing methods, there are still some challenges and possible future directions. We analyze several directions that are significant for future EEG-based emotion recognition.
Temporal fully-connected graph represents a fully-connected temporal graph across time slices. It aims to solve the problem of incomplete temporal dependency in current EEG-based emotion recognition. The temporal dependency of emotional EEG inferred by existing GNNs lacks the correlation between different channels across time slices. For Time series graph mentioned in Section 5.3, although the relationship between time slices can be established by the Temporal graph or encoder, this relationship is limited between the same electrodes across time slices, such as $v_{i}^{t-1}$ and $v_{i}^{t}$, while relations between different electrodes across time slice are ignored, such as $v_{i}^{t-1}$ and $v_{j}^{t}$. This hetero-electrode relationship across time slice corresponds to a delayed response in the emotional state of the brain, i.e., there is asynchrony in the interaction between brain regions. The existence of delays between specific brain regions and surrounding brain regions for brain region activity elicited by emotional excitation implies that this asynchronous relationship contains information directly related to emotion. Therefore, a more complete temporal dependency constructed by the Temporal fully-connected graph is expected to improve the accuracy of emotion recognition.
Graph condensation aims to compress a large and complex graph structure into a simple and compact representation and retain essential emotion-related information. Both coarsening and sparsifying are promising techniques for future EEG-based emotion recognition. Emotional EEG tends to move towards high resolution. Current GNNs in EEG-based emotion recognition utilize fully connected graphs, where nodes usually have an excessive number of neighbors, which provides redundant information and even noise. As the number of electrodes increases, this defect becomes more significant. Therefore, graph condensation techniques can not only improve training efficiency but also weaken the influence of emotionally irrelevant information. The activity of brain regions in emotional states is focused, which implies that the information of specific brain regions is sufficient to identify specific emotions, while most of the connections are emotionally irrelevant. In practice, GNNs in EEG-based emotion recognition are trained to preserve localization. Existing GNNs in this field generally need to run the training process at least once to find the optimal graph condensation, so if the optimal graph is identified earlier, e.g., by training for only a few epochs, then this will significantly accelerate the running process of GNNs to build emotional representation.
@@
Dynamic graph refers to a graph whose structure changes dynamically over time. It differs from dynamically updated GNNs in EEG-based emotion recognition, meaning the graph structure can be updated automatically as training proceeds. This belongs to model-level dynamic, while the dynamic graph is graph structure-level dynamic. Given dynamic graph $\textbf{G}(\textbf{V},\textbf{E})$, $\textbf{V}=\{(v,t_{s},t_{e})\}$ and $\textbf{E}=\{(e,t_{s},t_{e})\}$. $t_{s}$ and $t_{e}$ are timestamps when edges and nodes appear and disappear. Existing methods reason about the temporal dependency of emotional EEG by cropping time slices and constructing spatial graphs independently. The former indicates that the graph can be updated, and the latter indicates that what the graph represents is the dynamic temporal dependency in emotional EEG. Although temporal graphs or encoders can be used to build temporal dependencies between slices, such dependencies do not incorporate hetero-electrode relations across time slices, i.e., changes of edges across time. In contrast, the dynamic dependency depicted by the dynamic graphs encompasses changes in the edges and, therefore, corresponds to the complete time dependency of the emotional EEG.
@@ 7 Conclusion
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
{credits}
7.0.2 Acknowledgements
This study was supported in part by the Hubei Provincial Natural Science Foundation of China under Grant 2024AFB932, and in part by the CCF-NSFOCUS ’Kunpeng’ Research Fund.